OpenStudy (anonymous):

In sunlight, a vertical stick 9 ft tall casts a shadow 7 ft long. At the same time a nearby tree casts a shadow 28 ft long. How tall is the tree? Round to the nearest tenth

OpenStudy (anonymous):

you need to do proportion

OpenStudy (anonymous):

how would you set that up

OpenStudy (anonymous):

9/7 = 28/x then cross multiply and solve for x

OpenStudy (anonymous):

so 21.8

OpenStudy (anonymous):

wait no, 9/7 = x/28

OpenStudy (anonymous):

sorry i didnt read completely.

OpenStudy (anonymous):

so its 36

OpenStudy (anonymous):

yes